# SpecDec++: Boosting Speculative Decoding via Adaptive Candidate Lengths

### v2

@@ 3.2 SpecDec++
Given a threshold $h$, we end the current round of speculation when the predicted probability that there exists one rejected token exceeds $h$
$\pi(s_{k})={\mathrm{stop}}\Leftrightarrow\widehat{\mathbb{P}}(\exists 1\leq i\leq k,\text{ such that }Y_{i}\text{ is rejected}\mid x_{\text{prefix}})>h,$
@@
3.3 Training Dataset and Objective
Let $\mathcal{D}_{\text{prompt}}$ be the prompt distribution. For each $x_{\text{prompt}}$ in $\mathcal{D}_{\text{prompt}}$, we generate a response $(X_{1},\dots,X_{N})$ using the target model. Next, we feed the prompt and the response into the draft model to get $q(\cdot\mid x_{\text{prompt}},X_{1},\dots,X_{i-1})$ for every $i$. We sample a $Y_{i}$ from the distribution and calculate the conditional acceptance probability $\mathbb{P}_{i}=\min\Big{(}1,\frac{p(Y_{i}|x_{\text{prompt}},X_{1},\dots,X_{i-1})}{q(Y_{i}|x_{\text{prompt}},X_{1},\dots,X_{i-1})}\Big{)}$ for each token, which will be the training target.
We construct the response sequence $(Z_{1},\dots,Z_{N})$ by randomly taking $r\%$ tokens from $(X_{1},\dots,X_{N})$ and the remaining tokens from $(Y_{1},\dots,Y_{N})$, borrowing the random masking idea from BERT [14]. We only compute losses for the tokens from $(Y_{1},\dots,Y_{N})$.
We note that there will be distribution shifts between $(x_{\text{prefix}},Y_{1},\dots,Y_{k})$, the sequence encountered during the inference process, and $(x_{\text{prefix}},Z_{1},\dots,Z_{k})$, the sequence encountered during training process. The distribution shift may cause certain biases in the prediction head, e.g., over-confident about the acceptance. Furthermore, as in the typical setting of speculative decoding where the draft model and the target model align reasonably well, there will be class imbalance issues in the training dataset, where most of the training examples will have $\mathbb{P}_{i}$ close to $1$.
To accommodate the issues above, we train the prediction head using a weighted binary cross-entropy (BCE) loss, taken over the tokens $Z_{i}$’s stemming from $Y_{i}$’s:
@@ 4.1 Experimental Setups
Datasets and Model Pairs. We adopt three datasets in our experiments: Alpaca [39], HumanEval [9], GSM8K [13]. We only use prompts of the datasets and do not use responses. In the experiments, we use llama-2-chat models [42]. We choose to use llama-2-chat 7B as the draft model and llama-2-chat 70B as the target model. To reduce memory consumption, we use the bfloat16 format for the models.
@@
Baseline Method. We compare SpecDec++ with the naive speculative decoding algorithm where the number of the candidate tokens $K$ is fixed as a hyperparameter. We tune $K$ in $\{2,4,6,8,10,12,14\}$.
Due to space limits, additional experimental setup is deferred to Appendix C.1.
@@ 4.2 Forward Time Analysis
First, we verify the correctness of Equation (2.1) and determine the forward time of the draft model $t_{\text{draft}}$ and the target model $t_{\text{target}}$ under our specific setting. We collect all the $(N_{\text{draft}},N_{\text{target}},T_{\text{total}})$ tuples from generations using speculative decoding (either the baseline version or SpecDec++) and perform a linear regression to determine the coefficients. We also determine the standalone inference time when using only the draft model or the target model with linear regression. The linear regressions fit well with all $R^{2}\geq 0.98$ and the results are summarized in Table 2. Additionally, we visualize $t_{\text{draft}}$ and $t_{\text{target}}$ across the three settings in Figure 3.
Figure 3: The forward time of the draft model (llama-2-chat-7B) and the target model (llama-2-chat-70B) under different settings. For each setting, we perform linear regression to calculate the forward times and then average them across different datasets. The additional cost of the acceptance prediction head is negligible compared to the systematic error and the random noise of the environment. Full results are deferred to Table 2.
The additional cost of the acceptance prediction head is negligible, as we find that the average $t_{\text{draft}}$ in SpecDec++ setting is smaller than the average $t_{\text{draft}}$ in baseline SpecDec setting by $0.0004s$, which is likely caused by random noise of the environment, as the standard deviation between difference datasets around $0.0006s$. Therefore, for both the baseline speculative decoding setting and SpecDec++ setting, we choose $(t_{\text{draft}},t_{\text{target}})=(0.0234,0.112)$, which is the average between the two cases. We use Equation (2.3) to calculate the theoretical throughputs (tokens per second), which match the noisier empirical throughputs well with relative error $\leq 6.2\%$ for all prompts.
In the standalone setting where only the draft model or the target model is used, we see significant decreases in both $t_{\text{draft}}$ and $t_{\text{target}}$, which indicates that speculative decoding induces minor additional communication overhead. We use $(t_{\text{draft}},t_{\text{target}})=(0.0207,0.108)$ for the stand-alone setting. The average throughput for the target model is $9.26$ tokens/second.
4.3 Performances
@@
Figure 4: The average verification rate $N_{\text{target}}/N$ and the average discard rate $N_{\text{discarded}}/N$ for the baseline SpecDec under different candidate lengths and SpecDec++ with different acceptance prediction heads and stopping thresholds. SpecDec++ has better Pareto frontiers than SpecDec on both the in-distribution dataset Alpaca and the two out-of-distribution datasets HumanEval and GSM8K.
@@
Discussions. As the distribution shift of the OOD datasets will influence the accuracies and the calibrations of the acceptance prediction heads, a natural question to ask is whether the optimal performances for different datasets are achieved with different acceptance prediction heads and stopping thresholds. Empirically, we confirm that this is indeed the case. Nevertheless, we find that using the acceptance prediction trained with $w_{\text{rej}}=6$ and network depth $D=3$ and the stopping threshold $h=0.7$ achieves over 99.3% of the best tokens per second across the three datasets (2.03x for Alpaca, 2.21x for HumanEval, and 2.26x for GSM8K). Additional ablation studies on how the hyperparameters $(w_{\text{rej}},D,h)$ influence the final tokens per second can be found in Appendix C.3.
5 Related Work
Speculative decoding. Since the proposal of speculative decoding, people have been improving the algorithm from different perspectives. Our work is complementary to the works that improve speculative decoding by (1) making the draft model align better with the target model [54, 2, 29], (2) building smaller draft models or merging draft models into the target model (e.g. early-exiting) [30, 28, 49, 5, 51, 31, 12], and (3) building a heirachical system of speculative decoding [34, 37]. Our work is not directly appliable to the methods that do not have the concept of an auto-regressive draft model [35, 27, 6, 7] and the retrieval-based methods [17, 52, 47, 16]. See Appendix B for a detailed discussion on related work about speculative decoding, token trees, and diffusion language models.
Candidate length selection. Leviathan et al. [24] make the i.i.d. assumption on the acceptance probabilities of the candidate tokens and theoretically derive the optimal choice of $K$. Besides, Liu et al. [28] and Kim et al. [20] adopt a simple heuristic that ends the speculation if the confidence of the current draft token distribution falls below a threshold. Xu et al. [46] uses the cumulative product of the confidences and extends to the token tree version. In comparison, our work systematically studies the candidate length selection within the MDP framework and uses the cumulative product of our trained prediction head to determine the end of the speculation.
6 Conclusion
In this work, we study the determination of the candidate lengths for speculative decoding. We formulate the problem as a Markov Decision Process and provide a theorem that gives a sufficient condition to stop the current speculation. Motivated by the theoretical result, we propose SpecDec++ to adaptively select the candidate length with a trained acceptance prediction head. We demonstrate significant speedups over baselines and our method can be seamlessly integrated with other improvements.
@@ Acknowledgments
We would like to thank Tianle Cai, Kaifeng Lyu, Zhuoming Chen, and Beidi Chen for the helpful feedback on this paper.
@@ Appendix A Limitations
Our theoretical result contains a problem-specific constant $\Delta$ which is hard to analyze theoretically or estimate empirically. Nevertheless, the choice of the stopping threshold $h$ can be determined through hyperparameter search; see Appendix C.3. As is the case with all speculative decoding algorithms, our method relies on the implicit assumption that the draft model and the target model align well. For a weak draft model, the acceptance prediction head may perform badly.
@@ Appendix B Additional Related Work
Large language models are mostly based on Transformer architectures [43] that auto-regressively predict the probability of the next token given its predecessors. One bottleneck of the inference speed lies in the fact that auto-regressive decoding is an inherently non-parallelizable sequential operation: the probabilities of future tokens depend on the current token and there is no trivial way to skip the current token when predicting future tokens. Therefore, the inference time of auto-regressive decoding scales linearly with the number of the generated tokens.
However, the time of a forward pass to compute the log probabilities of the tokens through transformers is nearly constant for batched sequences with different lengths within a proper range, thanks to the increasingly powerful parallel computing units [32, 43, 8, 24].
Therefore, to overcome the bottleneck of the auto-regressive decoding, one can find a fast way to generate $K$ tokens, which often increases FLOPs, and the ask the target model to verify and correct the candidates [35, 8, 24]; see a comprehensive survey [45]. For those methods to work, we assume that we have enough computational resources (e.g. CUDA memories) to support the increased concurrency. Nevertheless, in the long-context generation regime, the memory issue becomes prominent, which requires additional KV-cache management techniques such as compression or retrieval [26, 37].
@@
The performance of speculative decoding depends on how well the draft model aligns with the target model, and how fast the draft model is compared to the target model. People have been improving speculative decoding in two aspects: (1) making the draft model align better with the target model via distillation [54, 2] and online learning [29]; and (2) making the token generation faster and cheaper, e.g. training multiple smaller draft models from stratch  [30].
In addition, the candidate tokens can be generated without a separate draft model [35, 27, 15, 6], such as building additional modules that predict the next $k$ tokens (Medusa heads [7], RNN heads [51], soft tokens [31]), early-exiting methods that reuse the intermediate representations of the target model [28, 49, 5], and retrieval-based methods that involve constructing an $n$-gram datastore and using retrieval to generate candidates [17, 52, 47, 16].
Those techniques can be combined, resulting in a heirachical system [34, 52, 37].
@@
Paralleling across the batch dimension via token trees is another direction to increase throughputs [30, 46, 36]. For greedy decoding, token tree generation and verification are studied in [7]. For the stochastic sampling setting, REST [17] proposes a straightforward approach: keeping the token paths that coincide with the stochastic tokens given by the target model. There are also researches extending the stochastic speculative decoding to the token tree setting, which often needs to adjust the drafting and verification probabilities to ensure unbiasedness, e.g. MCSD [48], Recursive SD [19], Sequoia [11], EAGLE [27], SpecTR [38].
@@
One important problem to study is how to construct and prune the token tree to maximize throughputs and avoid heavy communication overheads, which is studied in [11, 53]. Our work can serve as a starting point towards the problem, as the candidate length $K$ can be viewed as the depth of a token tree with only one branch.
@@
Diffusion language models. Diffusion language models either in the discrete space (see D3PM [4] and its follow-ups) or in the embedding space (see Diffusion-LM [25] and its follow-ups) are non-autoregressive language models, whose generation time can scale sub-linearly with the sequence length. BERT-type encoder-only models and auto-regressive decoder-only models can be also viewed as diffusion model, with mask prediction and next-token prediction being the denoising operation [4]. Viewing next-token prediction as Jacobi iteration [33] and denoising operation is a powerful idea and it leads to subsequent work such as lookahead decoding [16] and consistency LLMs [21].
Appendix C Additional Experimental Results
C.1 Additional Experimental Setups

### v3

@@ 4.3 SpecDec++ Algorithm
If we have such a head $f_{\theta}$, we propose to end the current round of speculation when the predicted probability that there exists one rejected token exceeds a predefined threshold $h$
$\pi(s_{k})={\texttt{stop}}\Leftrightarrow\widehat{\mathbb{P}}(\exists 1\leq i\leq k,\text{ such that }Y_{i}\text{ is rejected}\mid x_{\text{prefix}})>h,$
@@
4.4 Training Dataset Construction and Learning Objective
In this subsection, we focus on how to efficiently train an acceptance prediction head for SpecDec++. For each $x_{\text{prompt}}$ in the prompt set $\mathcal{D}_{\text{prompt}}$, we first generate a target response $(X_{1},\dots,X_{N})$ using the target model. Next, we feed the prompt and the response into the draft model to get $q(\cdot\mid x_{\text{prompt}},X_{1},\dots,X_{i-1})$ for every $i$. We sample a draft candidate $Y_{i}$ from the distribution and calculate the conditional acceptance probability $\mathbb{P}_{i}=\min\Big{(}1,\frac{p(Y_{i}|x_{\text{prompt}},X_{1},\dots,X_{i-1})}{q(Y_{i}|x_{\text{prompt}},X_{1},\dots,X_{i-1})}\Big{)}$ for each token, which will be the training target for the acceptance prediction head.
Token Mixing. Ideally, the input to the acceptance prediction head should be $(x_{\text{prompt}},X_{1},\dots,X_{i-1},Y_{i})$. However, this naive construction is training-inefficient as only the final token $Y_{i}$ receives a training signal. To overcome this, we propose a token mixing strategy, borrowing the random masking idea from BERT (Devlin et al., 2019): we randomly take $r\%$ tokens from $(X_{1},\dots,X_{N})$ and the remaining tokens from $(Y_{1},\dots,Y_{N})$ to construct the response sequence, denoted by $(Z_{1},\dots,Z_{N})$. The losses are only computed for the tokens from $(Y_{1},\dots,Y_{N})$. In this way, we trade the quality of the input sequences for training efficiency, as more tokens will receive training signals per forward pass of the model.
Weighted Binary Cross-Entropy (BCE) Loss.
In the typical setting of speculative decoding where the draft model and the target model align reasonably well, there will be class imbalance issues in the training dataset, where most of the training examples will have $\mathbb{P}_{i}$ close to $1$. To accommodate the issues above, we train the prediction head using a weighted binary cross-entropy (BCE) loss, taken over the tokens $Z_{i}$’s stemming from $Y_{i}$’s. In summary, our final loss function is
@@ 5.1 Experimental Setups
Datasets and Model Pairs. We adopt three datasets in our experiments: Alpaca (Taori et al., 2023), HumanEval (Chen et al., 2021), GSM8K (Cobbe et al., 2021). We only use prompts of the datasets and do not use responses. In the experiments, we solely focus on llama-2-chat models (Touvron et al., 2023b), while in Appendix D.4, we also provide additional experimental results on Gemma models (Team et al., 2024a; b). We choose to use llama-2-chat 7B as the draft model and llama-2-chat 70B as the target model. To reduce memory consumption, we use the bfloat16 format for the models.
@@
Baseline Method. We compare SpecDec++ with the simplest implementation of the speculative decoding algorithm where the number of the candidate tokens $K$ is fixed as a hyperparameter. We tune $K$ in $\{2,4,6,8,10,12,14\}$.
We report two metrics: (1) discard rate $N_{\text{discarded}}/N$, which is the average number of discarded tokens per one generated token, and (2) verification rate $N_{\text{target}}/N$, which is the average number of the forward calls of the target model per one generated token.
Due to space limits, additional experimental setup is deferred to Appendix D.1.
@@ 5.2 Forward Time Analysis
First, we verify the correctness of Equation (3.1) and determine the forward time of the draft model $t_{\text{draft}}$ and the target model $t_{\text{target}}$ under our specific setting. We collect all the $(N_{\text{draft}},N_{\text{target}},T_{\text{total}})$ tuples from generations using speculative decoding (either the baseline version or SpecDec++) and perform a linear regression to determine the coefficients. We also determine the standalone inference time when using only the draft model or the target model with linear regression. The linear regressions fit well with all $R^{2}\geq 0.98$ and the results are summarized in Appendix D.2, Table 2.
The additional cost of the acceptance prediction head is negligible, as we find that the average $t_{\text{draft}}$ in SpecDec++ setting is smaller than the average $t_{\text{draft}}$ in baseline SpecDec setting by $0.0004s$, which is likely caused by random noise of the environment. Therefore, for both the baseline speculative decoding setting and SpecDec++ setting, we choose $(t_{\text{draft}},t_{\text{target}})=(0.0234,0.112)$, which is the average between the two cases. For the stand-alone setting, we have $(t_{\text{draft}},t_{\text{target}})=(0.0207,0.108)$, indicating that the average throughput for the target model without speculative decoding is $9.26$ tokens/second.
5.3 Performance
@@
Figure 3: The average verification rates $N_{\text{target}}/N$ and the average discard rates $N_{\text{discarded}}/N$ for SpecDec with different candidate lengths and SpecDec++ with different acceptance prediction heads and stopping thresholds. SpecDec++ has better Pareto frontiers than SpecDec on both the in-distribution dataset Alpaca and the two out-of-distribution datasets HumanEval and GSM8K.
@@
Discussions. As the distribution shift of the OOD datasets will influence the accuracies and the calibrations of the acceptance prediction heads, a natural question to ask is whether the optimal performances for different datasets are achieved with different acceptance prediction heads and stopping thresholds. Empirically, we confirm that this is indeed the case. Nevertheless, we find that using the acceptance prediction trained with $w_{\text{rej}}=6$ and network depth $D=3$ and the stopping threshold $h=0.7$ achieves over 99.3% of the best tokens per second across the three datasets (2.03x for Alpaca, 2.21x for HumanEval, and 2.26x for GSM8K). Additional ablation studies on how the hyperparameters $(w_{\text{rej}},D,h)$ influence the final tokens per second can be found in Appendix D.3.
6 Conclusion and Discussion
In this paper, we study the determination of the candidate lengths for speculative decoding. We formulate the problem as a Markov Decision Process and provide a theorem that gives a sufficient condition to stop the current speculation. Motivated by the theoretical result, we propose SpecDec++ to adaptively select the candidate length with a trained acceptance prediction head. We demonstrate significant algorithmic speedups over the naive SpecDec baselines. Our paper focuses on the algorithmic aspects of speculative decoding with few assumptions on the system/hardware level implementations. Therefore, our method can be seamlessly integrated with other architectural or system-level improvements.
Theoretically, the acceleration achieved via SpecDec++ depends on the learnability of the conditional acceptance probability (Eqn. 4.1), which, in turn, depends on the amount of correlation between the degree of alignment between the draft distribution and the target distribution and the existing context. While it is hard to develop intuition into when such correlation is strong for a given pair of models, our empirical results demonstrate that the correlation indeed exists and can be extracted by training the acceptance prediction head, which ultimately leads to acceleration of speculative decoding. Meanwhile, the effectivenesses of many heuristic methods (Liu et al., 2024; Kim et al., 2024; Xu et al., 2023; Mamou et al., 2024) independently support the hypothesis that such correlation can be adequate for many pairs of models.
Our final remark is on the efficacy of speculative decoding in high-throughput scenarios. The conventional wisdom suggests that speculative decoding only improves latency and may hurt throughput when batch sizes are large. However, as pointed out by Sadhukhan et al. (2025), in the prevalent long-context settings, KV cache loading becomes the main bottleneck of inference, and this memory bottleneck cannot be amortized by increasing batch sizes. Speculative decoding becomes an effective method to achieve speedup in this long-context scenario, even when the batch sizes are large. We leave for future work the adaptation of our analysis and technique to the large-batchsize, long-context settings.
@@ Acknowledgments
We acknowledge Tianle Cai, Kaifeng Lyu, Zhuoming Chen, and Beidi Chen for the helpful feedback and discussion. Kaixuan Huang acknowledges the support of Google PhD Fellowship. Mengdi Wang acknowledges support by NSF grants DMS-1953686, IIS-2107304, and ONR grant 1006977. The research is also supported by Princeton Language and Intelligence (PLI) Compute Cluster.
@@ Appendix A Limitations
Our theoretical result contains a problem-specific constant $\Delta$ which is hard to analyze theoretically or estimate empirically. Nevertheless, the choice of the stopping threshold $h$ can be determined through hyperparameter search; see Appendix D.3. As is the case with all speculative decoding algorithms, our method relies on the implicit assumption that the draft model and the target model align well. For a weak draft model, the acceptance prediction head may perform badly.
@@ Appendix B Additional Related Work
Large language models are mostly based on Transformer architectures (Vaswani et al., 2017) that auto-regressively predict the probability of the next token given its predecessors. One bottleneck of the inference speed lies in the fact that auto-regressive decoding is an inherently non-parallelizable sequential operation: the probabilities of future tokens depend on the current token and there is no trivial way to skip the current token when predicting future tokens. Therefore, the inference time of auto-regressive decoding scales linearly with the number of the generated tokens.
However, the time of a forward pass to compute the log probabilities of the tokens through transformers is nearly constant for batched sequences with different lengths within a proper range, thanks to the increasingly powerful parallel computing units (Pope et al., 2023; Vaswani et al., 2017; Chen et al., 2023a; Leviathan et al., 2023).
Therefore, to overcome the bottleneck of the auto-regressive decoding, one can find a fast way to generate $K$ tokens, which often increases FLOPs, and the ask the target model to verify and correct the candidates (Stern et al., 2018; Chen et al., 2023a; Leviathan et al., 2023); see a comprehensive survey (Xia et al., 2024). For those methods to work, we assume that we have enough computational resources (e.g. CUDA memories) to support the increased concurrency. Nevertheless, in the long-context generation regime, the memory issue becomes prominent, which requires additional KV-cache management techniques such as compression or retrieval (Li et al., 2024a; Sun et al., 2024a).
@@
The performance of speculative decoding depends on how well the draft model aligns with the target model, and how fast the draft model is compared to the target model. People have been improving speculative decoding in two aspects: (1) making the draft model align better with the target model via distillation (Zhou et al., 2024; Agarwal et al., 2024) and online learning (Liu et al., 2023); and (2) making the token generation faster and cheaper, e.g. training multiple smaller draft models from stratch  (Miao et al., 2023).
In addition, the candidate tokens can be generated without a separate draft model (Stern et al., 2018; Li et al., 2024b; Du et al., 2024; Bhendawade et al., 2024), such as building additional modules that predict the next $k$ tokens (Medusa heads (Cai et al., 2024), RNN heads (Zhang et al., 2024), soft tokens (Monea et al., 2023)), early-exiting methods that reuse the intermediate representations of the target model (Liu et al., 2024; Yang et al., 2023b; Bae et al., 2023), and retrieval-based methods that involve constructing an $n$-gram datastore and using retrieval to generate candidates (He et al., 2023; Zhao et al., 2024; Yang et al., 2023a; Fu et al., 2024).
Those techniques can be combined, resulting in a heirachical system (Spector & Re, 2023; Zhao et al., 2024; Sun et al., 2024a).
@@
Paralleling across the batch dimension via token trees is another direction to increase throughputs (Miao et al., 2023; Xu et al., 2023; Su et al., 2023). For greedy decoding, token tree generation and verification are studied in (Cai et al., 2024). For the stochastic sampling setting, REST (He et al., 2023) proposes a straightforward approach: keeping the token paths that coincide with the stochastic tokens given by the target model. There are also researches extending the stochastic speculative decoding to the token tree setting, which often needs to adjust the drafting and verification probabilities to ensure unbiasedness, e.g. MCSD (Yang et al., 2024), Recursive SD (Jeon et al., 2024), Sequoia (Chen et al., 2024), EAGLE (Li et al., 2024b), SpecTR (Sun et al., 2024b).
@@
One important problem to study is how to construct and prune the token tree to maximize throughputs and avoid heavy communication overheads, which is studied in (Chen et al., 2024; Zhong et al., 2024). Our work can serve as a starting point towards the problem, as the candidate length $K$ can be viewed as the depth of a token tree with only one branch.
@@
Diffusion language models. Diffusion language models either in the discrete space (see D3PM (Austin et al., 2021) and its follow-ups) or in the embedding space (see Diffusion-LM (Li et al., 2022) and its follow-ups) are non-autoregressive language models, whose generation time can scale sub-linearly with the sequence length. BERT-type encoder-only models and auto-regressive decoder-only models can be also viewed as diffusion model, with mask prediction and next-token prediction being the denoising operation (Austin et al., 2021). Viewing next-token prediction as Jacobi iteration (Santilli et al., 2023) and denoising operation is a powerful idea and it leads to subsequent work such as lookahead decoding (Fu et al., 2024) and consistency LLMs (Kou et al., 2024).
B.1 Discussion: Sub-optimality of Heuristic Methods
In this section, we discuss the potential sub-optimality of several training-free heuristic methods for determining the candidate lengths. For example, it may be tempting to use the entropy of the draft model $\mathcal{H}(q(\cdot\mid x_{\text{prefix}},Y_{1},\dots,Y_{i-1}))$ or merely the likelihood of the sampled draft candidate token $q(Y_{i}\mid x_{\text{prefix}},Y_{1},\dots,Y_{i-1}))$ as a surrogate for the acceptance probability, and choose to stop the current round of speculation when these indicators fall under a threshold. However, we point out that this type of training-free heuristics confuses the inherent uncertainty of the draft distribution with the alignment between draft and target distributions, which is fundamentally flawed.
First of all, the theoretical acceptance probability is $\min\Big{(}1,\frac{p(Y_{i}|x_{\text{prefix}},Y_{1},\dots,Y_{i-1})}{q(Y_{i}|x_{\text{prefix}},Y_{1},\dots,Y_{i-1})}\Big{)}$, which depends on the target distribution $p$. Therefore, heuristic methods that do not incorporate the information of the target model will be sub-optimal.
Furthermore, when the draft model and the target model align well, the acceptance probability will be high regardless of the entropy of the current draft candidate token or the likelihood of the sampled $Y_{i}$. For the extreme case when the draft model and the target model aligns perfectly well i.e., $p(Y_{i}|x_{\text{prefix}},Y_{1},\dots,Y_{i-1})=q(Y_{i}|x_{\text{prefix}},Y_{1},\dots,Y_{i-1})$, by the rejection sampling scheme of speculative decoding, the sampled candidate tokens are guaranteed to be accepted, and the optimal candidate lengths will be infinite. Using either entropy-based methods or likelihood-based methods will stop the speculation earlier and results in sub-optimal performance.
Appendix C Additional Background on Speculative Decoding Algorithm
Rejection Sampling. The algorithmic foundation of the Speculative Decoding algorithm lies in rejection sampling. Specifically, if we want to sample from a target discrete distribution $p(x)$, we first sample $x$ from a draft distribution $q(x)$. We accept the sample $x$ with probability $\min(1,\frac{p(x)}{q(x)})$; otherwise we replace it with a sample from the modified distribution $\mathrm{Norm}[(p-q)_{+}]$, where $z_{+}=\max(z,0)$ is the positive part of $z$ and $\mathrm{Norm}[f]=\frac{f(\cdot)}{\sum_{x}f(x)}$ normalizes a function $f$ to make it a proper probability distribution. The proof of the unbiasedness of rejection sampling can be found in Chen et al. (2023a).
Speculative Decoding. Speculative decoding extends to the auto-regressive generation scenarios by chaining $K$ rejection sampling procedures together. The full algorithm is provided in Algorithm 2.
Algorithm 2 Speculative Decoding (Chen et al., 2023a; Leviathan et al., 2023)
  for $i=1$ to $K$  do
     Compute $q_{i}=q(\cdot\mid x_{\text{prefix}},y_{1},\dots,y_{i-1})$.
     Sample $y_{i}\sim q_{i}$.
  end for
   Sample $r_{1},\dots,r_{K}$ with $r_{i}\sim\mathrm{Unif}[0,1]$, $i=1,\dots,K$.
  if $n<K$ then
  else
     Sample $y^{\prime}$ from $p_{K+1}$
  end if
  Return $x_{\text{prefix}},y_{1},\dots,y_{n},y^{\prime}$
Appendix D Additional Experimental Results
D.1 Experimental Setups